# Massively Parallel Proof-Number Search for Impartial Games and Beyond

### v1

@@ Methods
We implemented SPOTS, a PNS-PDFPN-based solver for combinatorial games, augmented with Grundy numbers to reduce game trees of impartial games effectively. From now on, we use PNS and DFPN to refer to their Grundy-enhanced variants. We evaluate the performance of SPOTS on Sprouts, an impartial game well-suited for Proof-Number Search due to its highly unbalanced game trees, large branching factors, and the fact that it often naturally decomposes into independent subpositions. To incorporate Sprouts into SPOTS, we implemented the string-based position representation from applegate1991computer. Our experiments were carried out on the 29-spot and 47-spot positions, as these were the largest positions for which computations were completed in a reasonable time. The experimental results are reported as the mean $\pm$ standard error of the mean, based on three independent measurements.
@@
We also compare our PNS with the PNS implemented by lemoine2015nimber in their solver GLOP. Unlike our method, GLOP operates on a proper tree without transpositions and discards all visited nodes in the solved subtrees to reduce memory usage. Along with a roughly three-times faster state representation, our design choices lead to a roughly 100-fold reduction in computation time compared to GLOP while maintaining a comparable maximum tree size.
@@
The parallel PPN2 search by saffidine2012parallel corresponds to PNS-PNS using PNS at both levels. However, after each job is completed, the second-level PNS search tree is discarded, losing the entire worker state. In contrast, our PNS-DFPN workers retain their transposition tables, completely preserving their state throughout the whole computation. Figure 4 shows how the retention and sharing of the worker state impact scaling efficiency. Although PNS-PNS scales well to up to 64 cores, worker state resets hamper its performance so significantly that it is outperformed by sequential DFPN($10^{6}$). However, when the key results are retained, specifically the computed Grundy numbers, the speedup improves significantly. Further performance gains can be achieved by PNS-DFPN preserving computed proof and disproof numbers. However, this has a smaller effect than retaining the Grundy numbers. Finally, sharing Grundy numbers among workers yields a further significant performance gain.
@@
We evaluate the impact of the PNS-PDFPN parameters on the scaling efficiency, measured relative to the runtime of DFPN($5\cdot 10^{7}$), serving as a baseline for comparing various parallel Proof-Number Search algorithms. We compare PNS-PDFPN against PDFPN by kaneko2010parallel and the PNS-DFPN scheme employed by schaeffer2007checkers to solve Checkers, which we consider representative state-of-the-art approaches for shared- and distributed-memory systems, respectively.
In Figure 5, we incrementally add parameters until we arrive at the full PNS-PDFPN with all the improvements. For each setup, we report the speedup achieved using the best-performing parameter setting, where PNS-DFPN is already optimized for iterations and updates, as proposed by saffidine2012parallel. Without GN synchronization, PNS-DFPN scales very poorly. Enabling synchronization improves scalability up to 512 cores, but then the performance degrades due to significant synchronization overhead. This issue is overcome by grouping workers, reducing the amount of shared results, and thus the synchronization overhead. Further scalability is achieved by introducing second-level parallelization, which ultimately leads to a speedup of $208.67\pm 9.17$ without relying on any domain knowledge, which exceeds even the best speedup of $21.35\pm 2.40$ achieved by PDFPN using shared memory.
We observe a notable improvement in scaling efficiency after incorporating the child-ordering heuristic proposed by lemoine2015nimber to break ties during the selection of the next node $w$. Adding the heuristic does not enhance the overall performance if applied to optimally tuned parameters without the heuristic, as shown in the upper part of Table 1. However, the heuristic becomes beneficial when the parameters are tuned specifically for its use; see the bottom part of Table 1. The heuristic then enables more effective utilization of second-level parallelism in longer-running jobs by providing better guidance to branches that are more likely to yield shorter proofs. With this modest domain knowledge, the final speedup of PNS-PDFPN reaches $332.97\pm 26.8$. Further parameter analysis and hardware specification are given in Appendix.
Figure 5: Impact of GN synchronization (GN Sync), grouping (Gr), second-level parallelization (Th), and the child-ordering heuristic (Hr) on the final speedup of PNS-PDFPN relative to DFPN($5\cdot 10^{7}$), which computed for $293\pm 41.7$ minutes. Measured on the 47-spot position.
@@
With SPOTS, we determined 32 new outcomes of $n$-spot positions using the sequential version and 10 more with its parallel counterpart, nearly doubling the number of previously known results; see Appendix. We verified the new proofs using the verification software by lemoine2015nimber. All computed outcomes agree with the Sprouts Conjecture, which thus remains open. Compared to the previous state-of-the-art solver GLOP, our parallel version of SPOTS achieves a speedup of four orders of magnitude on 1024 cores. We base this estimate on the combined improvements of the sequential SPOTS over GLOP and the parallel SPOTS over the sequential version. To illustrate the speedup, SPOTS solves the $39$-spot position, one of the most complex solved by GLOP, in just 7.7 minutes using 1024 cores. The outcomes determined by parallel SPOTS are much more complex than those previously computed. The most difficult position was solved in 24 days on 512 cores, and its proof size (the number of stored Grundy numbers) exceeds prior ones by a factor of 1,000.
@@
Table 1: Best-performing values of iterations (It) and threads (Th) for PNS-PDFPN without (top) and with (bottom) the heuristic. Grouping set to its maximum and updates to 1000 are optimal across all experiments.
@@ Appendix A Acknowledgments
All authors were supported by the grant no. 25-18031S of the Czech Science Foundation (GAČR). M. Schmid was also supported by the Charles University project UNCE 24/SCI/008. T. Čížek was supported by the Charles University Grant Agency (GAUK) project no. 326525. The authors thank EquiLibre Technologies, Inc. for providing computational resources. Additional computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic. Special thanks to Neil Burch for valuable comments and insights.
@@ Appendix B Sprouts Game
Sprouts was designed by British mathematicians John Horton Conway and Michael Paterson in 1967 with the intention to create a simple-to-play but difficult-to-analyze game (berlekamp2003winning; roberts2015genius).
Since the early beginnings of Sprouts, researchers have tried to determine the outcomes of $n$-spot positions for as many values of $n$ as possible. Conway himself found the outcomes for $n\leq 3$. Mollison later proved the outcomes for $n$ equal to 4 and 5, as well as for $n=6$, whose proof by hand took 49 pages (gardner1967mathematical).
applegate1991computer created the first computer solver for Sprouts, which proved the outcomes for $n\leq 11$, and led the authors to formulate the Sprouts Conjecture. Their solver used Alpha-Beta Pruning with the string representation of positions on which all subsequent solvers, including ours, were built. In 2006, Josh Purinton created a computer program, called AuntBeast, that played casual games with players from the World Game Of Sprouts Association and was able to play a perfect game up to 14 spots; thus, newly solving the game for $n=12,13,14$.
The previous state-of-the-art solver GLOP created by lemoine2015nimber solved all $n$-spot positions for $n\leq 32$ in 2007. In 2011, they extended the results for $n\leq 44$ plus three more values to $n=53$; see Table 2. After hearing about these achievements made by Lemoine and Viennot, Conway, one of the authors of Sprouts, reacted with the following disbelieving words (roberts2015genius):
@@
The critical improvement by Lemoine and Viennot compared to the previous solver by applegate1991computer stands on the utilization of Grundy numbers to analyze independent subpositions separately and on human interventions to navigate the search outside of unpromising parts of a game tree. Their algorithm originally relied on Alpha-Beta Pruning with a heuristic for ordering child-node exploration. Later, Lemoine and Viennot replaced Alpha-Beta with a basic variant of Proof-Number Search, enabling them to achieve several new results.
@@ 21 W G 8e4 42 L G 1e6 63 84 105
Table 2: Solved $n$-spot Sprouts positions. Each entry in the $n$th row indicates the solver (C = Conway, M = Mollison, A = applegate1991computer, P = AuntBeast (Purinton 2006), G = GLOP (lemoine2015nimber), S = SPOTS), the outcome (W = win, L = loss), and size of the proof measured by the number of computed Grundy numbers. Lighter and darker G-entries indicate positions solved by GLOP in 2007 and in 2010–2011, respectively. Lighter and darker S-entries indicate positions solved by the sequential and parallel versions of our SPOTS solver, respectively.
@@
All the newly computed outcomes agree with the Sprouts conjecture, which thus remains open. Additionally, we confirm that the computed Grundy numbers of $n$-spot positions follow the extended Sprouts conjecture of lemoine2015nimber, which states that the Grundy number of the $n$-spot position with $n=0,1,\text{or }2\pmod{6}$ equals to 0, and to 1 otherwise.
@@
After a month of computations on this cluster with 480 cores allocated to workers, we derived proofs of Sprouts positions that are approximately 1,000 times larger than the most complicated ones derived by the current state-of-the-art solver GLOP (lemoine2015nimber; lemoine2011phdthesis). We computed around $10^{9}$ Grundy numbers in total, where one Grundy number corresponds to approximately 100 node expansions.
Notably, the results were achieved without fully utilizing the potential of our solver, as parameters were not always optimally tuned, and some workers were intermittently terminated during computation due to spot machine preemption. The parallel SPOTS computations were executed on 480 CPU cores, primarily using the following parameter settings: workers = 480, iterations = 10,000, updates = 1000, grouping = 3, and threads = 10.
@@
The supplementary materials include the source code of our parallel solver, SPOTS, for the game of Sprouts. We also provide a database of Grundy numbers as a certificate for the newly solved positions. Due to size constraints, we attach only the database corresponding to positions solved by the sequential solver. The full database of positions solved by the parallel solver exceeds the supplementary material size limit and will be made publicly available on GitHub after the review process, along with the source code. A link will be included in the final version.
All computation-demanding parts of the solver, such as the representation of a state and search algorithms, are implemented in C++20 for performance reasons. The first-level work distribution is implemented in Python $3.10$ using Ray $2.47.0$, a framework for distributed systems. The master and the workers are coordinated using Ray, while they utilize the core functions implemented in C++, exposed to Python using pybind11 $2.10$. The whole project is built using CMake $3.22$ with the compiler GCC $11.4$. The project targets Linux operating systems and is tested on Ubuntu 22.04 LTS.
The SPOTS solver can compute the outcome of any given Sprouts position. Positions are provided as input using the string representation introduced by applegate1991computer and later used by lemoine2015nimber. In particular, the $n$-spot position is encoded as 0*$n$.
@@
To estimate the mean and variance of game tree complexity for Sprouts, we sampled 1,000 random games by selecting each child uniformly at random at each node along a path $\mathcal{P}$ from the root to a terminal node. Let $b_{i}$ denote the number of children of the $i$th node along $\mathcal{P}$. We computed $\prod_{i=1}^{k}b_{i}$ for each path, where $k$ is the number of internal nodes in the path. The final estimate was then obtained by averaging these values across all the samples, with the standard deviation reported as a measure of variance.
To estimate the complexity of the Sprouts tree with Grundy numbers, we used a similar method based on randomly sampled games, but with an important modification: decomposable nodes must be treated separately, as their total complexity is the sum of the complexities of subtrees of all their children. The estimation is therefore recursive and instead of the path $\mathcal{P}$ in the tree, we obtain a randomly sampled subtree $\mathcal{T}$. As in our description of DFPN with GN, we first simplify the Sprouts tree by replacing Grundy nodes with atomic nodes, resulting in a tree that includes only atomic and decomposable nodes. The subtree $\mathcal{T}$ is then obtained by starting at the root, selecting a single child at every encountered atomic node uniformly at random, and by selecting all children of each encountered decomposable node. After sampling $\mathcal{T}$, we initially set the complexity $C(t)$ of each terminal $t$ in $\mathcal{T}$ to be 1. To compute the complexity $C(v)$ of the Sprouts subtree of an internal node $v$ of $\mathcal{T}$, we distinguish two cases. If $v$ is atomic, then we let $C(v)=b(v)\cdot C(c)$, where $c$ is the child of $v$ in $\mathcal{T}$ and $b(v)$ is the number of children of $v$ in the Sprouts tree. For decomposable $v$, we let $C(v)=\sum_{c}C(c)$, where the sum is taken over all the children $c$ of $v$ in the Sprouts tree. The complexity estimate of the Sprouts tree with Grundy numbers is then the average of the values $C(r)$ of the root $r$ taken over 1,000 samples of $\mathcal{T}$, with the range between the 1st and 99th percentiles reported as the measure of variance. During the sampling, we need to estimate the sizes of the Grundy numbers of the nodes. However, these values are unknown. Using the average values of the Grundy numbers in the computed proof trees of $n$-spot positions, we estimate the number of relevant children of Grundy nodes to determine their value.

### v2

@@ Methods
We implemented SPOTS, a PNS-PDFPN-based solver for combinatorial games, augmented with Grundy numbers to reduce game trees of impartial games effectively. From now on, we use PNS and DFPN to refer to their Grundy-enhanced variants. We evaluate the performance of SPOTS on Sprouts, an impartial game well-suited for Proof-Number Search due to its highly unbalanced game trees, large branching factors, and the fact that it often naturally decomposes into independent subpositions. To incorporate Sprouts into SPOTS, we implemented the string-based position representation from Applegate et al. (1991). Our experiments were carried out on the 29-spot and 47-spot positions, which are the largest that allow each trial to complete within 24 hours. The experimental results are reported as the mean $\pm$ standard error of the mean, based on three independent measurements.
@@
We also compare our PNS with the PNS implemented by Lemoine and Viennot (2015) in their solver GLOP. Unlike our method, GLOP operates on a proper tree without transpositions and discards all visited nodes in the solved subtrees to reduce memory usage. Along with a roughly three-times faster state representation, our design choices lead to a roughly 100-fold reduction in computation time compared to GLOP while maintaining a comparable maximum tree size.
@@
The parallel PPN2 search by Saffidine et al. (2012) corresponds to PNS-PNS using PNS at both levels. However, after each job is completed, the second-level PNS search tree is discarded, losing the entire worker state. In contrast, our PNS-DFPN workers retain their transposition tables, completely preserving their state throughout the whole computation. Figure 4 shows how the retention and sharing of the worker state impact scaling efficiency. Although PNS-PNS scales well to up to 64 cores, worker state resets hamper its performance so significantly that it is outperformed by sequential DFPN($10^{6}$). However, when the key results are retained, specifically the computed Grundy numbers, the speedup improves significantly. Further performance gains can be achieved by PNS-DFPN preserving computed proof and disproof numbers. However, this has a smaller effect than retaining the Grundy numbers. Finally, sharing Grundy numbers among workers yields a further significant performance gain.
@@
We evaluate the impact of the PNS-PDFPN parameters on the scaling efficiency, measured relative to the runtime of DFPN($5\cdot 10^{7}$), serving as a baseline for comparing various parallel Proof-Number Search algorithms. We compare PNS-PDFPN against PDFPN by Kaneko (2010) and the PNS-DFPN scheme employed by Schaeffer et al. (2007) to solve Checkers, which, according to our measurements, represent state-of-the-art approaches for shared- and distributed-memory systems, respectively.
In Figure 5, we incrementally add parameters until we arrive at the full PNS-PDFPN with all the improvements. For each setup, we report the speedup achieved using the best-performing parameter setting, where PNS-DFPN is already optimized for iterations and updates, as proposed by Saffidine et al. (2012). Without GN synchronization, PNS-DFPN scales very poorly. Enabling synchronization improves scalability up to 512 cores, but then the performance degrades due to significant synchronization overhead. This issue is overcome by grouping workers, reducing the amount of shared results, and thus the synchronization overhead. Further scalability is achieved by introducing second-level parallelization, which ultimately leads to a speedup of $208.67\pm 9.17$ without relying on any domain knowledge, which exceeds even the best speedup of $21.35\pm 2.40$ achieved by PDFPN using shared memory.
We observe a notable improvement in scaling efficiency after incorporating the child-ordering heuristic proposed by Lemoine and Viennot (2015) to break ties during the selection of the next node $w$. Adding the heuristic does not enhance the overall performance if applied to optimally tuned parameters without the heuristic, as shown in the upper part of Table 1. However, the heuristic becomes beneficial when the parameters are tuned specifically for its use; see the bottom part of Table 1. The heuristic then enables more effective utilization of second-level parallelism in longer-running jobs by providing better guidance to branches that are more likely to yield shorter proofs. With this modest domain knowledge, the final speedup of PNS-PDFPN reaches $332.97\pm 26.8$. Further parameter analysis and hardware specification are given in Appendix.
Figure 5: Impact of GN synchronization (GN Sync), grouping (Gr), second-level parallelization (Th), and the child-ordering heuristic (Hr) on the final speedup of PNS-PDFPN relative to DFPN($5\cdot 10^{7}$), which finished in $293\pm 41.7$ minutes. Measured on the 47-spot position.
@@
With SPOTS, we determined 32 new outcomes of $n$-spot positions using the sequential version and 10 more with its parallel counterpart, nearly doubling the number of previously known results; see Appendix. We verified the new proofs using the verification software by Lemoine and Viennot (2015). All computed outcomes agree with the Sprouts Conjecture, which thus remains open. Compared to the previous state-of-the-art solver GLOP, our parallel version of SPOTS achieves a speedup of four orders of magnitude on 1024 cores. We base this estimate on the combined improvements of the sequential SPOTS over GLOP and the parallel SPOTS over the sequential version. To illustrate the speedup, SPOTS solves the $39$-spot position, one of the most complex solved by GLOP, in just 7.7 minutes using 1024 cores. The outcomes determined by parallel SPOTS are much more complex than those previously computed. The most difficult position was solved in 24 days on 512 cores, and its proof size (the number of stored Grundy numbers) exceeds prior ones by a factor of 1,000.
@@
Table 1: Best-performing values of iterations (It) and threads (Th) for PNS-PDFPN without (top) and with (bottom) the heuristic. Grouping set to its maximum and updates to 1,000 are optimal across all experiments.
@@ Appendix A Acknowledgments
All authors were supported by the grant no. 25-18031S of the Czech Science Foundation (GAČR). T. Čížek was supported by the Charles University Grant Agency (GAUK) project no. 326525. M. Schmid was also supported by the Charles University project UNCE 24/SCI/008. The authors thank EquiLibre Technologies, Inc. for providing computational resources. Additional computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic. Special thanks to Neil Burch for valuable comments and insights.
@@ Appendix B Sprouts Game
Sprouts was designed by British mathematicians John Horton Conway and Michael Paterson in 1967 with the intention to create a simple-to-play but difficult-to-analyze game (Berlekamp et al. 2003; Roberts 2015).
Since the early beginnings of Sprouts, researchers have tried to determine the outcomes of $n$-spot positions for as many values of $n$ as possible. Conway himself found the outcomes for $n\leq 3$. Mollison later proved the outcomes for $n$ equal to 4 and 5, as well as for $n=6$, whose proof by hand took 49 pages (Gardner 1967).
Applegate et al. (1991) created the first computer solver for Sprouts, which proved the outcomes for $n\leq 11$, and led the authors to formulate the Sprouts Conjecture. Their solver used Alpha-Beta Pruning with the string representation of positions on which all subsequent solvers, including ours, were built. In 2006, Josh Purinton created a computer program, called AuntBeast, that played casual games with players from the World Game Of Sprouts Association and was able to play a perfect game up to 14 spots; thus, newly solving the game for $n=12,13,14$.
The previous state-of-the-art solver GLOP, created by Lemoine and Viennot (2015), solved all $n$-spot positions for $n\leq 32$ in 2007. In 2011, they extended the results for $n\leq 44$ plus three more values to $n=53$; see Table 2. After hearing about these achievements made by Lemoine and Viennot, Conway, one of the authors of Sprouts, reacted with the following disbelieving words (Roberts 2015):
@@
The critical improvement by Lemoine and Viennot compared to the previous solver by Applegate et al. (1991) stands on the utilization of Grundy numbers to analyze independent subpositions separately and on human interventions to navigate the search outside of unpromising parts of a game tree. Their algorithm originally relied on Alpha-Beta Pruning with a heuristic for ordering child-node exploration. Later, Lemoine and Viennot replaced Alpha-Beta with a basic variant of Proof-Number Search, enabling them to achieve several new results.
@@ 21 W G 8e4 42 L G 1e6 63 84 105
Table 2: Solved $n$-spot Sprouts positions. Each entry in the $n$th row indicates the solver (C = Conway, M = Mollison, A = Applegate et al. (1991), P = AuntBeast (Purinton 2006), G = GLOP (Lemoine and Viennot 2015), S = SPOTS), the outcome (W = win, L = loss), and size of the proof measured by the number of computed Grundy numbers. Lighter and darker G-entries indicate positions solved by GLOP in 2007 and in 2010–2011, respectively. Lighter and darker S-entries indicate positions solved by the sequential and parallel versions of our SPOTS solver, respectively.
@@
All the newly computed outcomes agree with the Sprouts conjecture, which thus remains open. Additionally, we confirm that the computed Grundy numbers of $n$-spot positions follow the extended Sprouts conjecture of Lemoine and Viennot (2015), which states that the Grundy number of the $n$-spot position with $n=0,1,\text{or }2\pmod{6}$ equals to 0, and to 1 otherwise.
@@
After a month of computations on this cluster with 480 cores allocated to workers, we derived proofs of Sprouts positions that are approximately 1,000 times larger than the most complicated ones derived by the current state-of-the-art solver GLOP (Lemoine and Viennot 2015; Lemoine 2011). We computed around $10^{9}$ Grundy numbers in total, where one Grundy number corresponds to approximately 100 node expansions.
Notably, the results were achieved without fully utilizing the potential of our solver, as parameters were not always optimally tuned, and some workers were intermittently terminated during computation due to spot machine preemption. The parallel SPOTS computations were executed on 480 CPU cores, primarily using the following parameter settings: workers = 48, iterations = 10,000, updates = 1,000, grouping = 3, and threads = 10.
@@
The project’s GitHub repository contains the source code for our parallel Sprouts solver, SPOTS. It also provides links to the Grundy number databases, which serve as certificates for newly solved positions.
All computation-demanding parts of the solver, such as the representation of a state and search algorithms, are implemented in C++20 for performance reasons. The first-level work distribution is implemented in Python $3.10$ using Ray $2.47.0$, a framework for distributed systems. The master and the workers are coordinated using Ray, while they utilize the core functions implemented in C++, exposed to Python using pybind11 $2.10$. The whole project is built using CMake $3.22$ with the compiler GCC $11.4$.
The SPOTS solver can compute the outcome of any given Sprouts position. Positions are provided as input using the string representation introduced by Applegate et al. (1991) and later used by Lemoine and Viennot (2015). In particular, the $n$-spot position is encoded as 0*$n$.
@@
To estimate the mean and variance of game tree complexity for Sprouts, we sampled 1,000 random games by selecting each child uniformly at random at each node along a path $\mathcal{P}$ from the root to a terminal node. Let $b_{i}$ denote the number of children of the $i$th node along $\mathcal{P}$. We computed $\prod_{i=1}^{k}b_{i}$ for each path, where $k$ is the number of internal nodes in the path. The final estimate was then obtained by averaging these values across all the samples, with the range between the 1st and 99th percentiles as the measure of variance.
To estimate the complexity of the Sprouts tree with Grundy numbers, we used a similar method based on randomly sampled games, but with an important modification: decomposable nodes must be treated separately, as their total complexity is the sum of the complexities of subtrees of all their children. The estimation is therefore recursive, and instead of the path $\mathcal{P}$ in the tree, we obtain a randomly sampled subtree $\mathcal{T}$. As in our description of DFPN with GN, we first simplify the Sprouts tree by replacing Grundy nodes with atomic nodes, resulting in a tree that includes only atomic and decomposable nodes. The subtree $\mathcal{T}$ is then obtained by starting at the root, selecting a single child at every encountered atomic node uniformly at random, and by selecting all children of each encountered decomposable node. After sampling $\mathcal{T}$, we initially set the complexity $C(t)$ of each terminal $t$ in $\mathcal{T}$ to be 1. To compute the complexity $C(v)$ of the Sprouts subtree of an internal node $v$ of $\mathcal{T}$, we distinguish two cases. If $v$ is atomic, then we let $C(v)=b(v)\cdot C(c)$, where $c$ is the child of $v$ in $\mathcal{T}$ and $b(v)$ is the number of children of $v$ in the Sprouts tree. For decomposable $v$, we let $C(v)=\sum_{c}C(c)$, where the sum is taken over all the children $c$ of $v$ in the Sprouts tree. The complexity estimate of the Sprouts tree with Grundy numbers is then the average of the values $C(r)$ of the root $r$ taken over 1,000 samples of $\mathcal{T}$, with the range between the 1st and 99th percentiles reported as the measure of variance. During the sampling, we need to estimate the sizes of the Grundy numbers of the nodes. However, these values are unknown. Using the average values of the Grundy numbers in the computed proof trees of $n$-spot positions, we estimate the number of relevant children of Grundy nodes to determine their value.